# Chemical Language Model Linker: blending text and molecules with modular adapters

### v1

@@ 3.7 Baseline methods
T5 The T5 model comes from the MolT5 paper (Edwards et al., 2022). It is trained on the ChEBI-20 dataset.
@@
TGM-DLM TGM-DLM (Gong et al., 2024) uses the text embedding to guide a diffusion language model to generate molecules. It employs a two-phase diffusion process, first generating a molecule based on the text description followed by correcting invalid SMILES strings.
@@
T5 encoder+MolXPT MLP This model concatenates the T5 text and MolXPT molecule embeddings and passes them through a two-layer multilayer perceptron (MLP). It serves as an ablation for ChemLML to demonstrate the value of the full ChemLML model over its individual components.
@@ 3.8 Training and evaluation
We evaluate the model’s performance on text-based molecule generation. For MolT5, Text+ChemT5, MolXPT, SciBERT, and different scales of MolGen, we use their checkpoints on HuggingFace. For the Text+ChemT5 encoder, we add the prompt "Write in SMILES the described molecule:", which is used in the corresponding task in Text+ChemT5.
For ChemLML, we use SciBERT, the encoder from Text+ChemT5 and different scales of Galactica as the text encoders and combine them with MolGPT, MolXPT, MolGen, and MolGen-7B as the molecule decoders. The weights of molecule decoder are always frozen. The weights of the adapter are trainable. The weights of the text encoder can be either trainable or frozen. We run experiments on two datasets, ChEBI-20 and PubChem.
For ChEBI-20, we report the results from the MolT5, Text+ChemT5, MolXPT, and TGM-DLM papers. T5 is included in the MolT5’s paper.
MolGPT can not be finetuned in ChEBI-20 because it is pretrained on MOSES dataset so its tokenizer can not deal with chirality information. We use MolGPT to compare SMILES- and SELFIES-based models. First, we use RDKit to remove stereochemistry information from all the molecules. Then, we remove all the molecules that fail to be tokenized by MolGPT’s tokenizer. Lastly, we transform all the SMILES to SELFIES. This reduces the ChEBI-20 training set from 26,407 to 15,899 instances, validation set from 3,301 to 1,961 instances, and test set from 3,300 to 2,032 instances. We build the MolGPT model with 12 layers, 8 attention heads, and 256 as the embedding dimension. The only differences between MolGPT SMILES and SELFIES models are the output dimensions, depending on the vocabulary lists of each method.
We use the Noam Optimizer with 4,000 warm up steps for training. For the sampling in the test set, we use multinomial sampling with random seed 42 for all methods.
@@ 3.9 Docking case study setting
We filtered the PubChem dataset to identify text descriptions that pertain to molecules that inhibit specific protein targets by initially retaining only descriptions with the string "inhibit". Then, we split all remaining instances into low (0.15 to 0.3), medium (0.3 to 0.5), and high (0.5 to 0.9) similarity based on the Morgan fingerprint similarity between the ground truth and generated molecule from the ChemLML (T5 encoder+MolGen) model. We selected one example per similarity bin (Table 7.4) as described in the Supplementary Methods.
After we generated the first molecule with the temperature set to 1 and random seed 42, we set the temperature to 1.5 and changed the random seed in order to sample different molecules. We canonicalized the generated SMILES with RDKit and removed duplicates. We repeated this process until we obtained 99 additional unique and valid SMILES, resulting in a total of 100 ChemLML-generated molecules per target. The SMILES were then used to build 3D conformers with OpenEye’s Omega2, and only a subset produced valid 3D structures for docking (Supplementary Methods). For ChemLML, we used the T5 encoder+MolXPT and T5 encoder+MolGen settings. We omit "encoder" for brevity in the docking experiments. We performed the same molecule generation process with the MolT5 and Text+ChemT5 baselines.
For each target protein, we docked the ground truth molecule, generated molecules, and control molecules from two types of background distributions: FDA-approved compounds and ChemLML T5+MolGen-generated molecules from randomly-sampled descriptions. A consensus docking score was computed for each molecule that combines results from four docking programs (Supplementary Methods). We hypothesized that the ground truth molecule and generated molecules resembling this ground truth molecule would score more favorably on their respective protein targets than background molecules.
3.10 Code and data availability
@@ 4 Results
We first compare baseline and ChemLML models trained on ChEBI-20 text descriptions on the ChEBI-20 test set (Table 1). The best performing ChemLML model is the T5 encoder finetune+MolXPT version. With 114M trainable parameters, this model achieves 0.727 in Morgan FTS. It is not surprising to find that models that include the T5 encoder work the best because the T5 encoder has been trained on multiple tasks. The ChemLML combination T5 encoder+MolXPT performs better than the baseline MolT5 even though MolT5 has 52 times more trainable parameters.
@@ 4.2 Molecule similarity and validity trade-off
There is a trade-off between generated and ground truth molecule similarity and the generated molecule validity (Table 1). T5 and MolT5 have the same model architecture. MolT5 has been pretrained on a large amount of data with both SMILES and natural language while T5 has not. At first glance, it is therefore surprising that T5 outperforms MolT5 in the similarity metrics. However, MolT5 generates 11 percentage points more valid molecules than T5, which means more molecules are counted in the similarity calculation. These molecules that are more difficult to generate may have lowered the overall similarity.
@@ 4.3 SMILES versus SELFIES representations
The choice of molecule representation, SMILES or SELFIES, is another factor that influences the similarity calculations. Skinnider (2024) recently showed that a molecule language model trained on SMILES had Fréchet ChemNet Distance (FCD) lower than 0.5, while one trained on SELFIES had FCD greater than 2 because invalid SMILES are low-likelihood samples and they improve performance on distribution-learning metrics like FCD. However, MolGen is trained on SELFIES, and there is not yet an equivalent model pretrained on SMILES that goes through multiple stages of pretraining. Therefore, we use MolGPT to compare SMILES and SELFIES directly.
@@
After MolGPT models are pretrained on the MOSES dataset, we train ChemLML T5 encoder+MolGPT models on the ChEBI-20 training dataset and evaluate the performance on the test set (Table 2). There is a large difference between the performance of SMILES and SELFIES versions of the same model. When we limit the molecules to the intersection of valid molecules between both methods to ensure the results are unaffected by the similarity and validity tradeoff we mention above, the SELFIES method still performs poorly. The SMILES method outperforms the SELFIES method by about 50% in the Morgan fingerprint similarity metric. Therefore the prior study’s conclusions about unconditional molecule generation (Skinnider, 2024) also hold for conditional molecule generation.
Based on SMILES substantially outperforming SELFIES in this analysis, we focus on comparing different pretrained models instead of comparing between SMILES- and SELFIES-based methods.
@@ 4.4 PubChem evaluation reveals pros and cons of LLMs
Because many of the PubChem molecule descriptions are overly ambiguous and generic, we prioritize the PubChem-filtered version of the dataset but also evaluate the PubChem-unfiltered dataset to assess how the filtering affects the results. On PubChem-filtered, LLMs without finetuning perform poorly, while the finetuned LLMs encoders show better results (Table 3). PubChem-filtered has a similar distribution as the ChEBI-20 dataset, so the ChemLML T5 encoder+MolXPT model still outperforms other methods that have a frozen text encoder. The ChemLML combination T5 encoder finetune+MolGen 7B does not perform well, potentially due to overfitting on the ChEBI-20 dataset. Another notable finding is that ChemLML generates a higher proportion of exactly matched molecules compared to the MolT5 and Text+ChemT5 baselines. We suggest that it is because MolGen and MolXPT have been pretrained on a large number of molecules, then ChemLML is able to retrieve those molecules accurately based on the text embedding.
@@
Table 3: Results of molecule generation on the PubChem-filtered test set. SciBERT is excluded because its encoder cannot handle inputs exceeding 512 tokens. The ChemLML models are grouped into those that finetune the text model and those that do not, and we bold the best model in each of those categories.
@@ 4.5 Docking case study
To demonstrate how ChemLML could be used in practice, we applied it to generate chemical inhibitors of protein drug targets acetylcholinesterase, inosine-5’-monophosphate dehydrogenase, and heat shock protein HSP 90-alpha (Table 7.4). We obtained the experimental structures of these proteins and used FRED (McGann, 2011), Gnina (McNutt et al., 2021), PLANTS (Korb et al., 2009), and rDock (Ruiz-Carmona et al., 2014) to dock the ground truth inhibitor (co-crystallized ligand) from PubChem, the ChemLML-generated candidate inhibitors, and candidate inhibitors generated by two baseline methods. In addition, we docked control molecules generated by ChemLML and an FDA-approved screening library that are not expected to bind the targets. We generated a single consensus docking score per compound (Figure 2).
Figure 2: Docking score distributions for the three protein targets in Table 7.4. Higher scores on the y-axis are better. Sample sizes for the docked compound sets are: FDA 2,956 for all targets, ChemLML_background (bg) 657 for all targets, ChemLML T5+MolGen 16/42/52, ChemLML T5+MolXPT 99/99/92, MolT5 97/98/96, and Text+ChemT5 83/96/86.
Docked ground truth and most generated molecules score poorly for the heat shock protein HSP 90-alpha, likely due to poor handling of macrocycles by docking programs (Table 7.4). Here, control compounds from the FDA and ChemLML background sets have median scores (0.00009 and 0.00043, respectively) similar to the ground truth molecule (0.00006) and the ChemLML T5 + MolGen generated ensemble median (0.00004). The ChemLML T5+MolXPT generated ensemble scores slightly higher with a median of 0.00115 but not substantially better than the random ChemLML background. Considering the poor docking score of the ground truth molecule, we consider the HSP 90-alpha target a poor system for comparative performance evaluation of the molecule generators. Also, we cannot draw conclusions about the quality of the 3D conformers built from the SMILES of the generated molecules.
@@
For inosine-5’-monophosphate dehydrogenase, the ground truth molecule and ChemLML T5+MolGen ensemble score markedly better (0.00488 and median of 0.00594) than the control FDA and ChemLML background sets (medians of 0.00006 and 0.00049). The ChemLML T5+MolXPT ensemble, however, does not show more favorable scores with a median score near the backgrounds (0.00040). The difference in median scores could be related to the difference in valid 3D conformations produced by each of these ChemLML variants.
@@
For acetylcholinesterase, the ChemLML T5+MolGen generated molecules have docking scores (median 0.00237) well above the background distributions for FDA and generated sets (medians of 0.00013 and 0.00064) but below the ground truth molecule (0.00873). The other generated ensembles—ChemLML T5+MolXPT, MolT5, and Text+ChemT5—do not perform better than the background sets. The first generated molecule from ChemLML T5+MolGen has a docking score (0.00482) above the background distributions, yet departs structurally from the ground truth molecule (Table 7.4), which is valuable in drug discovery. We also inspected whether the generated molecules satisfied the chemical structural properties in the text descriptions and their chemical structural diversity (Supplementary Results).

### v2

@@ 3.7 Baseline methods
T5 The T5 model comes from the MolT5 paper 28. It is trained on the ChEBI-20 dataset.
@@
TGM-DLM TGM-DLM 36 uses the text embedding to guide a diffusion language model to generate molecules. It employs a two-phase diffusion process, first generating a molecule based on the text description followed by correcting invalid SMILES strings.
@@
T5 encoder+MolXPT MLP We introduce this baseline model as an ablation for ChemLML to demonstrate the value of the full ChemLML model over its individual components. This model concatenates the T5 text and MolXPT molecule embeddings and passes them through a two-layer multilayer perceptron (MLP).
@@ 3.8 Training and evaluation
We evaluate the models’ performance on text-based molecule generation. For MolT5, Text+ChemT5, MolXPT, SciBERT, and different scales of MolGen, we use their checkpoints on HuggingFace. For the Text+ChemT5 encoder, we add the prompt “Write in SMILES the described molecule”, which is used in the corresponding task in Text+ChemT5.
For ChemLML, we use SciBERT, the encoder from Text+ChemT5, and different scales of Galactica as the text encoders and combine them with MolGPT, MolXPT, MolGen, and MolGen-7B as the molecule decoders. The weights of the molecule decoder are always frozen. The weights of the adapter are trainable. The weights of the text encoder can be either trainable or frozen. We run experiments on two datasets, ChEBI-20 and PubChem.
For ChEBI-20 baseline evaluations, we obtain the results from the MolT5, Text+ChemT5, MolXPT, and TGM-DLM publications. The T5 results are also obtained from the MolT5 publication.
MolGPT can not be finetuned on ChEBI-20 because it is pretrained on the MOSES dataset so its tokenizer does not interpret chirality information. We use MolGPT to compare SMILES- and SELFIES-based models. First, we use RDKit to remove stereochemical information from all of the molecules. Then, we remove the molecules that fail to be tokenized by MolGPT’s tokenizer. Lastly, we transform all the SMILES to SELFIES. This reduces the ChEBI-20 training set from 26,407 to 15,899 instances, validation set from 3,301 to 1,961 instances, and test set from 3,300 to 2,032 instances. We build the MolGPT model with 12 layers, 8 attention heads, and a 256-dimensional embedding. The only differences between the MolGPT SMILES and SELFIES models are the output dimensions, depending on the vocabulary lists of each method.
We use the Noam Optimizer with 4,000 warm-up steps for training. For molecule sampling, we use multinomial sampling with a random seed of 42 for all methods unless noted otherwise.
@@ 3.9 Docking case study setting
We filtered the PubChem dataset to identify text descriptions that pertain to molecules that inhibit specific protein targets by initially retaining only descriptions with the string “inhibit”. Then, we split all remaining instances into low (0.15 to 0.3), medium (0.3 to 0.5), and high (0.5 to 0.9) similarity based on the Morgan fingerprint similarity between the ground truth and generated molecule from the ChemLML(T5 encoder+MolGen) model. We selected multiple examples per similarity bin (Table 8.4) as described in the Supplementary Methods.
After we generated the first example molecule with the temperature set to 1 and random seed 42, we set the temperature to 1.5 and changed the random seed in order to sample different molecules. However, in some cases, the conditions specified in the conditional molecule generation task were overly restrictive. The model generated only a few unique molecules even when sampling 1,000 times. In these cases, we iteratively increased the temperature to encourage diversity and obtain the desired number of unique molecules. We generated up to 1,000 molecules per temperature. If the target number of unique molecules had not yet been obtained, we increased the temperature. The maximum temperature we reached with this procedure was 4.5. The iterative process and variability across prompts made it difficult to track the number of effective sampling iterations required to generate the target number of unique molecules. As a representative example, we report the number of duplicate samples in the permeability case study (Table 4).
We canonicalized the generated SMILES with RDKit and removed duplicates. We repeated the iterative generation process above until we obtained 99 additional unique and valid SMILES, resulting in a total of 100 ChemLML-generated molecules per target. The SMILES were then used to build 3D conformers with OpenEye’s Omega2, and only a subset produced valid 3D structures for docking (Supplementary Methods). For ChemLML, we used the T5 encoder+MolXPT and T5 encoder+MolGen settings. We omit “encoder” for brevity in the docking experiments. We performed the same molecule generation process with the MolT5 and Text+ChemT5 baselines.
For each of eight target proteins, we docked the ground truth molecule, generated molecules, and control molecules from two types of background distributions: FDA-approved compounds and ChemLML(T5+MolGen)-generated molecules from randomly-sampled descriptions. A consensus docking score was computed for each molecule that combines results from four docking programs (Supplementary Methods). We hypothesized that the ground truth molecule and generated molecules resembling this ground truth molecule would score more favorably on their respective protein targets than background molecules.
3.10 Permeability case study setting
We initially used the prompt “The molecule has high membrane permeability”. However, we found that MolT5 and Text+ChemT5 generated molecules containing protons such as “[H+].[H+].[H+].CN(C=O)C=O.CI” and nonsensical combinations of natural and chemical language like “CN (C)CI via minimal irritation on minimal water condition.CNC”. Although these can be parsed by RDKit, they do not represent chemically valid or meaningful molecules. Therefore, we modified the prompt to “The molecule is a drug-like compound with high passive membrane permeability. It contains no formal charge and avoids ionizable groups”. We used a script to remove invalid molecules, elemental substances, salt fragments, and other irrelevant side chains, retaining only the main structure. A generated result is considered valid only if it remains unchanged after this filtering process. We used a temperature of 2.0 and continued sampling by switching the seed until we obtained the desired number of molecules.
3.11 Data and Software Availability
@@ 4 Results
We first compare baseline and ChemLML models trained on ChEBI-20 text descriptions on the ChEBI-20 test set (Table 1). Some versions of the ChemLML models finetune the text encoder and some use the frozen text encoder. The best performing ChemLML model is the T5 encoder finetune+MolXPT version. With 114M trainable parameters, this model achieves 0.727 in Morgan FTS. It is not surprising to find that models that include the T5 encoder work the best because the T5 encoder has been trained on multiple tasks. The ChemLML combination T5 encoder+MolXPT performs better than the baseline MolT5 even though MolT5 has 52 times more trainable parameters.
@@ 4.2 Molecule similarity and validity trade-off
There is a trade-off between generated and ground truth molecule similarity and the generated molecule validity (Table 1). T5 and MolT5 have the same model architecture. MolT5 has been pretrained on a large amount of data with both SMILES and natural language while T5 has not. At first glance, it is therefore surprising that T5 outperforms MolT5 in the similarity metrics. However, MolT5 generates 11 percentage points more valid molecules than T5, which means more molecules are evaluated by the similarity calculation. These molecules that are more difficult to generate may have reduced the mean similarity.
@@ 4.3 SMILES versus SELFIES representations
The choice of molecule representation, SMILES or SELFIES, is another factor that influences the similarity calculations. Skinnider 60 recently showed that when training molecule language models on samples from ChEMBL, molecules generated from SMILES-based models matched the training set much better than SELFIES-based models. Invalid SMILES are low-likelihood samples, and filtering low-likelihood outputs improves performance on distribution-learning metrics. However, MolGen is trained on SELFIES, and there is not yet an equivalent model pretrained on SMILES that goes through multiple stages of pretraining. Therefore, we use MolGPT to compare SMILES and SELFIES directly.
@@
After MolGPT models are pretrained on the MOSES dataset, we train ChemLML T5 encoder+MolGPT models on the ChEBI-20 training dataset and evaluate the performance on the test set (Table 2). There is a large difference between the performance of SMILES and SELFIES versions of the same model. When we limit the molecules to the intersection of valid molecules between both methods to ensure that the results are unaffected by the similarity and validity trade-off we mention above, the SELFIES method still performs poorly. The SMILES method outperforms the SELFIES method by about 50% in the Morgan fingerprint similarity metric. Therefore, the conclusions of the prior study about unconditional molecule generation 60 also hold for conditional molecule generation.
Given that SMILES substantially outperform SELFIES in this analysis, we focus on comparing different pretrained models instead of comparing SMILES- and SELFIES-based methods.
@@ 4.4 PubChem evaluation reveals pros and cons of LLMs
We use molecule descriptions in PubChem as an additional molecule generation evaluation beyond ChEBI-20. Because many of the PubChem molecule descriptions are overly ambiguous and generic, we prioritize the PubChem-filtered version of the dataset but also evaluate the PubChem-unfiltered dataset to assess how filtering affects the results. On PubChem-filtered, LLMs without finetuning exhibit limited performance compared with T5 encoders that have been trained on relevant examples. After finetuning, LLM encoders show comparable results (Table 3). PubChem-filtered has a similar distribution as the ChEBI-20 dataset, so the ChemLML T5 encoder+MolXPT model still outperforms other methods that have a frozen text encoder. The ChemLML combination T5 encoder finetune+MolGen 7B does not perform well, potentially due to overfitting on the ChEBI-20 dataset. Another notable finding is that ChemLML generates a higher proportion of exactly matched molecules compared to the MolT5 and Text+ChemT5 baselines. We suggest that because MolGen and MolXPT have been pretrained on a large number of molecules, ChemLML is able to retrieve those molecules accurately based on the text embedding.
@@
Table 3: Results of molecule generation on the PubChem-filtered test set. SciBERT is excluded because its encoder cannot handle inputs exceeding 512 tokens. The ChemLML models are grouped into those that finetune the text model and those that do not. The highest value in each performance category is shown in bold.
@@ 4.5 Docking case study
To demonstrate how ChemLML could be used in practice, we applied it to generate chemical inhibitors of eight protein drug targets: acetylcholinesterase (AChE), inosine-5’-monophosphate dehydrogenase (IMPDH), heat shock protein HSP 90-alpha (HSP90AA1), SARS coronavirus main proteinase (Mpro), lysine-specific histone demethylase 1A (LSD1), DNA topoisomerase 2-beta (TOPIIB), angiotensin-converting enzyme (ACE), and mitogen-activated protein kinase kinase 1 (MAPKK1). Details of the targets and text prompts are shown in Table 8.4. We obtained the X-ray crystallographic structures of these proteins and used FRED 61, Gnina 62, PLANTS 63, and rDock 64 to dock the ground truth inhibitor (co-crystallized or known high-affinity ligand) from PubChem, the ChemLML-generated candidate inhibitors, and candidate inhibitors generated by two baseline methods. In addition, we docked control molecules generated by ChemLML and an FDA-approved screening library that are not expected to bind the targets. We generated a single consensus docking score per compound (Figure 2). Figures S2-S9 contain docking score distributions from the four individual docking programs. They demonstrate the benefits of the consensus docking approach. For example, on IMPDH, PLANTS and Gnina fail, but FRED, rDock, and the resulting consensus score succeed (Figure S3).
Figure 2: Higher scores on the y-axis are better. Sample sizes for the docked compound sets are displayed in Table 8.4. ECR: exponential consensus ranking
@@
In four of eight targets, the ChemLML(T5+MolGen) compounds dock with median scores exceeding those for their respective ground-truth ligand molecules. For IMPDH, the ground truth molecule, ribavirin monophosphate, and ChemLML(T5+MolGen) model score markedly better (0.00488 and median of 0.00594) than the control FDA and ChemLML background sets (medians of 0.00006 and 0.00049). ChemLML(T5+MolXPT), however, fails to improve over the backgrounds (0.00040). The difference in median scores between the ChemLML variants could be related to the difference in valid 3D conformations produced by each of these ChemLML variants. Median docking scores for ChemLML(T5+MolGen)-generated compounds for targets LSD1 (0.00249), ACE1 (0.00688), and MAPKK1 (0.00452) surpass those of their respective the ground truth compounds (0.00092, 0.00174, and 0.00445). In each case, the ground truth molecule and ChemLML(T5+MolGen) results show substantially better scoring than the control FDA and ChemLML background sets.
@@
In two of the eight targets, AChE and TOPIIB, the ChemLML(T5+MolGen)-generated compounds score better than control FDA and random ChemLML-generated sets but contain only a few compounds that exceed the ground truth molecules’ scores. For AChE, the ChemLML(T5+MolGen)-generated molecules have docking scores (median 0.00237) well above the background distributions for FDA and generated sets (medians of 0.00013 and 0.00064) but below the ground truth molecule (0.00873). The other generated compounds from ChemLML(T5+MolXPT), MolT5, and Text+ChemT5 do not score better than the background sets. The first generated molecule from ChemLML(T5+MolGen) has a docking score (0.00482) above the background distributions, yet departs structurally from the ground truth molecule (Table 8.4), which is valuable in drug discovery. We also inspected whether the generated molecules satisfied the chemical structural properties in the text descriptions and their chemical structural diversity (Supplementary Results). In the case of TOPIIB, the median docking scores for both ChemLML(T5+MolGen) (0.00052) and ChemLML(T5+MolXPT) (0.00124) both exceed those for the FDA (0.00007) and ChemLML-random (0.00011) background sets. However, the ground truth molecule (amsacrine–a DNA intercalator) docks more favorably (0.00723).
ChemLML-generated compounds fair poorly against two of the eight targets, HSP90AA1 and Mpro. Docked ground truth and most generated molecules score poorly for HSP90AA1, likely due to poor handling of macrocycles (Table 8.4) by docking programs. Here, control compounds from the FDA and ChemLML background sets have median scores (0.00009 and 0.00043, respectively) similar to the ground truth molecule (0.00006) and the ChemLML(T5+MolGen) median (0.00004). The ChemLML(T5+MolXPT) median scores slightly higher (0.00115) but not substantially better than the random ChemLML background. Considering the poor docking score of the ground truth molecule, we consider HSP90AA1 a fraught target for comparative performance evaluation of the molecule generators. Also, we cannot draw conclusions about the quality of the 3D conformers built from the SMILES of the generated molecules. In the case of Mpro, we see favorable scoring (0.00077) for the ground truth molecule (Savinin) but poor median scoring for both ChemLML models: T5+MolGen (0.00007) and T5+MolXPT (0.00006). The co-crystallized ligand for Mpro is a 5-mer peptide, which occupies a large substrate binding cavity, potentially conferring selectivity for larger (higher molecular weight) ligand compounds.
During molecule generation, we also found a limitation of the models jointly trained on the mixture of natural language and chemicals. When we increase the temperature to produce diverse molecules, the models collapse and generate natural language text mixed with SMILES strings. This phenomenon worsens as molecule descriptions become less structured. We discuss this further in our permeability experiment below.
4.6 Permeability case study
Assessing a chemical’s interactions with membrane transporters is an important part of drug development because those interactions can affect drug absorption, disposition, and excretion 65. In vitro membrane permeability is one chemical property assessed for this purpose, which can be measured with a MDR1-MDCK efflux ratio (ER) assay. The assay determines whether a chemical is a substrate of MDR1, also known as P-glycoprotein, which controls drug export and is expressed in the brain endothelia, intestine, kidney, and liver 65. A chemical is typically considered a substrate if the ER is $\geq 2$ 65. We assess whether ChemLML and existing text-based molecule generation models can generate membrane permeable compounds. Because the prompt (Methods) does not contain information about the MDR1 protein or the MDR1-MDCK ER assay, this case study serves as a more general test of the models’ ability to generate compounds with complex properties relevant for drug development.
We compare ChemLML(T5+MolXPT) and ChemLML(T5+MolGen) with the baselines MolT5 and Text+ChemT5. All models run until they produce 100 unique molecules that pass multiple validity filters, and we assess how many duplicate molecules are generated and the percentage of unique molecules that pass these filters (Table 4). Then, we use an existing MDR1-MDCK ER model 66 to predict the membrane permeability scores for these 100 generated molecules per model (Figure 3). Because ER $\geq 2$ typically indicates the molecule is a MDR1 substrate, the generated permeable molecules should have a predicted MDR1-MRCK ER score $<2$.
All models generally succeed in generating membrane permeability molecules, and the distributions of predicted MDR1-MDCK ER scores are similar for three of the models. ChemLML(T5+MolXPT) lags behind the other three models with a notable upward shift in the score distribution and some generated molecules that are predicted to be MDR1 substrates. A possible explanation is that MolXPT itself is fine-tuned on ChEBI-20. We only use its molecule generation capability, thus it may not generalize well on this out-of-distribution description that occurs infrequently in the ChEBI-20 text. In addition, we observe that ChemLML(T5+MolXPT) generates the highest number of duplicate molecules (Table 4), suggesting that it tends to converge on specific structures rather than generating diverse molecules. Overall, ChemLML(T5+MolGen) demonstrates the highest success rate in molecule generation with the fewest filtered molecules and effectively produces molecules that are predicted to be permeable, as specified by the prompt. By leveraging a single-modality molecular language decoder (MolGen), this ChemLML model avoids generating natural language tokens (like T5) while still effectively following natural language instructions.
Methods Sample Duplicate Invalid NL Salts SE Success rate (%)
MolT5 260 31 113 8 4 4 43.7
Text+ChemT5 984 28 318 395 101 42 10.6
ChemLML(T5+MolXPT) 624 519 3 1 1 0 95.2
ChemLML(T5+MolGen) 105 3 1 0 0 1 98.0
Table 4: Statistics of generated molecules for the permeability experiment. Each method generates molecules until it generates 100 successful molecules.
Sample: Number of total molecule generation sampling iterations
Duplicate: Number of duplicate molecules among the generated samples.
$Sample-Duplicate=Unique$
Success rate (%): The percentage of molecules that successfully pass all four filters below relative to the number of unique generated molecules. $100/Unique$
These four cases are considered as an unsuccessful generation:
Invalid: Number of generated molecules that could not be parsed by RDKit.
NL (natural language): Number of molecules containing natural language text or tokens outside the molecular representation space.
Salts: Number of molecules containing salt components.
SE (single element): Number of molecules composed of only a single chemical element.
$Unique$ equals the 100 successfully generated molecules plus the number of filtered molecules across these four categories.
Figure 3: Distribution of MDR1-MRCK ER scores of different models. MDR1-MRCK ER score < 2 indicates predicted permeability.